evaluate ∫ (x^e + e^x + e^e) dx

evaluate
∫ (x^e + e^x + e^e) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (x^e+e^x+e^e)dx
I = \int x^edx+\int e^xdx+\int e^edx
I = \frac{x^{e+1}}{e+1}+\int e^xdx+\int e^edx
I = \frac{x^{e+1}}{e+1}+e^x+\int e^edx
I = \frac{x^{e+1}}{e+1}+e^x+e^e.x + c